Explaination range in projectile motion.


1
Expert's answer
2021-04-18T19:33:34-0400

Range of a projectile motion refers to the total horizontal distance covered by the particle during the course of its motion (for simplicity in x direction).

In the above figure,

Velocity in x direction = "vcos\theta"

Total time of flight of projectile = "T"

Range of projectile, "R=vcos\theta\times T"